-- FeliceCusano - 31 May 2019

INAF queue observing, 2019 May 31- June 01

Observer: F. Cusano
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn (on call)
Instrument: LBC

At the sunset the sky is clear. We start with LBC. For a moth in LBC-B we switched to LUCI at 5:20. The OSA performed some test on the stability of LUCI

Summary

Night duration (12 degree el twilight) 3:23 - 11:11 UT = 7 h 48m

Observing time: 5h 51min (75%)

Weather time loss: 0h 0min (0%)

technical time loss 1h *57min (25%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
calibration sky flat LBC-B/R g; z     good counts
Annibali R27 LBC-B/R g; r 0.33 in g, 0.66 in r /1h(bino)   red images are good, blue have a spot moving in the fov (maybe a moth?) to be checked
telluric HIP87643 LUCI1-2 G210 J 0.15/0.15h   Casasola
Casasola mask1 LUCI1-2 G210 J 2.22/5h    
Casasola mask1 LUCI1-2 imaging J 0.17/0.5h    
telluric HIP100448 LUCI1-2 G200 zJspec 0.15/0.15h   Cresci
Cresci PSO J308.04 LUCI1-2 G200 zJspec 2.16/2h   completed
Cresci PSO J308.04 LUCI1-2 imag J /0.07h   completed
calibration flat and arc LUCI1-2 G200 zJspec     for Cresci
calibration flat and arc LUCI1-2 G210 J     for Casasola
calibration flat LUCI1-2 imag J      

Details

01:34 error intializing LBC, LBC B rotator error
  • 10:43 problem solved turning off and then on the LBC*

01:35 starting dark with LUCI
  • 10:36 mos error on luci2
  • 2:09 problem solved, starting dark for Moretti, Telluric_Moretti, Casasola, Telluric_Casasola
  • the luc1 dark are completed, only for 4 dark of Moretti are good for luci2

2:30 the TO opens the dome

2:48 starting sky flat g; z
  • some r and i also but counts are low

3:12 slewing to Annibali R27
  • 3:16 dohybrid
  • 3:32 copointing
  • 3:27 calibration ob UGCA276_calib, fwhm on images 0.8-1"
  • 3:31 starting science UGCA276_1, fwhm on images 0.9-1.1, some stars in the blue chanel are elongated we stop at 3:40 for dofpia
  • 3:41 dofpia
  • 3:45 sending again UGCA276_1, fwhm on images 0.7-0.9"
  • 4:10 dofpia
  • 4:14 sending UGCA276_2, fwhm on images 0.7-1.0", stars in the blue images are still bit elongated, but the seeing is good and the red channel looks ok, stop at 4:34 the star on blue images look still elongated
  • 4:34 dofpia
  • 4:40 sending again UGCA276_2, we take some images more, some of the previous were affected by elongation in the blue channel
  • 4:48 images looks ok now fhwm 0.7-0.8", but the first lbc-b image have a strange spot between ccd2 and ccd3, red image still ok
  • 4:57 stop integration the spot is still there, it is maybe a moth out of focus, the to and osa try to move the filter wheel
  • 5:09 starting again UGCA276_2, moth is still there
  • moth is still there we change instrument we switch to LUCI

5:21 switching to LUCI

5:48 Olga is performing dither test for the transformation of LUCI

6:42 slewing to Casasola program we start with the telluric
  • acq
  • 6:43 dimm seeing 0.9"
  • 6:44 mos error on LUCI2, 6:56 problem solved
  • 6:58 mos error again on LUCI2, 7:02 problem solved
  • 7:03 mos error again on LUCI2
  • 7:18 starting telluric acq luci1.20190601.0070-72; luci2.20190601.0055-57
  • telluric spec luci1.20190601.0073-76; luci2.20190601.0058-61
  • 7:24 moving to science field, filter J
  • acq science luci1.20190601.0077-80; luci2.20190601.0062-65
  • spec science luci1.20190601.0081-100; luci2.20190601.0066-85
  • 7:58 dimm seeing 1"
  • 8:38 dimm seein 1.2"
  • 8:48 fwhm on guider 1"
  • 9:05 imag J luci1.20190601.0101-105; luci2.20190601.0086-90

09:17 slewing to Cresci program, telluric HIP100448
  • acq tell luci1.20190601.0106-108; luci2.20190601.0091-95
  • 9:27 dimm seeing 0.8"
  • spec tell luci1.20190601.0109-112; luci2.20190601.0096-99
  • 9:39 moving to science
  • acq science luci1.20190601.0113-116; luci2.20190601.0100-103
  • spec science luci1.20190601.0117-142; luci2.20190601.0104-129
  • 10:00 dimm seeing 1.1"
  • 10:58 dimm seeing 0.9"
  • 11:13 sending imaging script J filter luci1.20190601.0143-150; luci2.20190601.0130-137

11:32 starting luci calibration
  • Cresci flat and arc
  • 11:35 luci2 mos error!
  • 11:43 problem solved
  • 11:51 starting flat and arc for Casasola
  • 11:52 mos error, solved at 12:08
  • 12:26 starting imaging flat in J
Topic revision: r7 - 07 Jun 2019, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback